OpenStudy (anonymous):

Part 1: Show the expanded form and simplification of . x^6/x^2Explain in your own words how you can simplify x^6/x^2 without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Would the expanded form be: \[ x^6x^{-2} \]? If so, then, we know: \[ x^ax^b=x^{a+b} \]Which means that: \[ x^6x^{-2}=x^{6-2}=x^4 \]

OpenStudy (anonymous):

i think thts how ecxpanded form goes thanx